Ali's age now is his dad's when Ali was 5 years old. When his dad reaches 71, Ali will be his dad's age now. What is the age for both of them now?
Answer by ankor@dixie-net.com(22740)

What is the age for both of them now?
:
let a = Ali's present age
let d = Dad's present age
:
let x = (d-a), the difference in their ages
:
"Ali's age now is his dad's when Ali was 5 years old."
a = x + 5
:
" When his dad reaches 71, Ali will be his dad's age now."
d = 71 - x
a + x = d
d = d, therefore
a + x = 71 - x
a = 71 - x - x
a = 71 - 2x
replace a with (x+5) from the first statement
x + 5 = 71 - 2x
x + 2x = 71 - 5
3x = 66
x = 66/3
x = 22 yrs is the difference in their ages
therefore
a = 22 + 5
a = 27 yrs is Ali's age
and
d = 71 - 22
d = 49 yrs is Dad's age, 22 yrs older than Ali